# Non-Hermiticity Induced Universal Anomalies in Kondo Conductance

### v1

@@ Abstract
Introduction.–Non-Hermitian physics has recently provided fresh insights into condensed matter systems. It leads to a wealth of novel quantum phenomena, including the non-Hermitian skin effect Zhao et al. (2025b); Yoshida et al. (2024); Song et al. (2019a); Okuma et al. (2020); Kawabata et al. (2023); Li et al. (2020b, a), non-Hermitian topology Ochkan et al. (2024); Yao and Wang (2018); Bergholtz et al. (2021); Yang et al. (2024); Song et al. (2019b); Hamanaka et al. (2024); Kawabata et al. (2019); Gong et al. (2018), and exceptional points Lee (2022); Ding et al. (2022); Tang et al. (2020); Miri and Alù (2019). These phenomena predominantly arise in non-interacting or weakly interacting systems, where quasi-particle descriptions are well-defined and applicable. However, for strongly correlated systems where the notion of quasi-particle fundamentally breaks down, the role of non-Hermiticity remains largely unexplored.
@@
In this letter, we uncover that the interplay between nFL and nH can generate anomalous Kondo conductance behaviors absent in all existing Kondo systems. To ground our theory in physical settings, we propose a practical setup that realizes the non-Hermitian multichannel Kondo (nHMCK) model (Eq.(1) below). The perturbative renormalization group (RG) analysis predicts both a weak-coupling and a strong-coupling phase, as shown in Fig.1. Around the weak-coupling FP, the impurity moment decouples from the conduction electrons even for antiferromagnetic Kondo couplings, leading to a universal low-temperature conductance, $\sim 1/\mathrm{ln}^{2}(T/T_{K})$. Although the conductance is similar to that of the underscreened Kondo models  Posazhennikova and Coleman (2005); Coleman and Pépin (2003); Mehta et al. (2005), it is driven by a completely different dissipation mechanism–the nH-induced decoupling. Beyond the weak-coupling phase, Bethe ansatz approach gives richer physics than that predicted by the perturbative RG analysis Kattel et al. (2025d, a), which is verified by our non-Hermitian numerical renormalization group (NRG) approach. Importantly, after imposing $\mathcal{PT}$ symmetry, it is found that the low-energy physics of the strong coupling phase behaves similarly with that of the Hermitian case, and is described by a boundary conformal field theory (BCFT). Remarkably, the fractionalized nature of the nFL, when combined with non-Hermiticity, causes an anomalous deviation in the Kondo conductance with raising the temperature (Fig.4 below). Such a nH-enriched anomaly stems from an emergent BCFT with non-Hermitian boundary operators, which is unique to systems exhibiting both nH and nFL. These results point to a novel class of transport phenomena unrecognized before, driven simultaneously by correlation and dissipation.
@@
Away from the weak-coupling FP, the situation is more subtle, as the perturbative analysis could breakdown. Along the red arrow in Fig.3(b), a critical curve (marked by yellow) occurs, separating the weak and strong non-Hermiticity region. Notably, for an intermediate non-Hermiticity region, Bethe ansatz studies on the single-channel case revealed an interesting Yu-Shiba-Rusinov (YSR)-like phase Kattel et al. (2025d). Here, we solve the Bethe ansatz equation for the multichannel model without $\mathcal{PT}$ symmetry, and find that the YSR-like phase still exists in this region (Sec.V of Ref. Yi et al. (2026).). In the weak non-Hermiticity region with small $\theta$, although the perturbative RG predicts the flow towards $(g_{r},g_{i})=(2/n,0)$, non-perturbative calculations are indispensable to fully understand the low-energy Kondo physics. Thus, we further utilize the non-Hermitian NRG Burke and Mitchell (2025) to solve Eq.(1), taking the $n=2$ case as an example. The details of our NRG method are presented in Ref. Yi et al. (2026).
As shown by Fig.3(e)(f), the real spectrum at large RG step coincides with that of the Hermitian case for small $\theta$, while significant deviations are observed for large $\theta$. Meanwhile, the impurity entropy saturates to $\mathrm{ln}2/2$ and $\mathrm{ln}2$ for small and large $\theta$, respectively (Fig.3(d)). These indicate that the system indeed flows to the overscreened (unscreened) phase under weak (strong) non-Hermiticity. For small $\theta$, the entropy decreases monotonously with temperature (blue curve in Fig.3(d)), displaying the standard Kondo regime behavior. For large $\theta$, the local moment phase is identified for both the high and low temperature regime, again validating the circling RG flow in Fig.3(b)  Nakagawa et al. (2018). Remarkably, for both large and intermediate $\theta$, the impurity entropy exhibits a non-monotonous evolution and can exceed the free local moment entropy, $\ln 2$, in an intermediate temperature range (red curve in Fig.3(d)). Such anomalous behaviors were also reported in superconductors with an impurity  Kattel et al. (2025b, c) . They reflect the fact that additional degrees of freedom emerge and contribute to the impurity entropy, providing clear evidence of the YSR phase predicted by Ref. Kattel et al. (2025d).
@@
Conclusion and discussion.– We have shown that non-Hermiticity can generate unconventional Kondo conductance behaviors. Particularly, when nFL meets nH, an anomalous upturn would occur, originating from an emergent BCFT with non-Hermitian boundary operators. The non-Hermitian LIO identified here is unique for Kondo systems that simultaneously display nH and nFL. For example, for the non-Hermitian Fermi liquid FP (which occurs for $n=1$) , the impurity is exactly screened, and the LIOs around the strong-coupling FP are simply four-fermion operators without imaginary couplings  [^3] The four-fermion operators with imaginary couplings break $\mathcal{PT}$, thus are prohibited around the FP.. Here, we emphasize that these BCFT results are applicable only to $\mathcal{PT}$ symmetric model, while for the asymmetric cases, our NRG calculations have confirmed the emerging YSR phase Kattel et al. (2025d, a), beyond BCFT descriptions. In addition, it is worthwhile to consider generalizations to other correlated systems exhibiting nFL properties, such as the Luttinger liquid Lee and Toner (1992); Fröjdh and Johannesson (1995); Furusaki and Nagaosa (1994); Han et al. (2023); Tang et al. (2024). Furthermore, we remark that multi-orbital cold-atom systems also provide a feasible platform to realize our predicted Kondo phenomena, as they can be engineered to host both localized and itinerant degrees of freedom  Gorshkov et al. (2010); Riegger et al. (2018). Designable dissipation and system-integrable monitoring equipments are also achievable in this platform  Syassen et al. (2008); Takasu et al. (2020); Bouganne et al. (2020); Zhao et al. (2025a, c); Vanhoecke and Schirò (2025). Last, our predicted anomalous conductance could serve as an additional signature of Majorana fermions. Given its strong-correlation and nonlocal origin, this signature may help distinguish Majorana modes from trivial bound states in experimental measurements.
@@ Acknowledgements.
W. Z. Y and Y. C contributed equally to this work. We acknowledge Wei-Qiang Chen for inspiring discussions. This work was supported by the National Natural Science Foundation of China (No.12322402, No.12274206, No.125B2076), the Scientific Research Innovation Capability Support Project for Young Faculty (SRICSPYF-ZY2025164), the National R&D Program of China (2024YFA1410500, 2022YFA1403601), the Innovation Program for Quantum Science and Technology (Grant No.2021ZD0302800), the Natural Science Foundation of Jiangsu Province (No.BK20233001), and the Fundamental Research Funds for the Central Universities (KG202501).

### v2

# arXiv:2602.13803v2[cond-mat.str-el] 02 Mar 2026

@@ Abstract
Introduction.–Non-Hermitian physics has recently provided fresh insights into condensed matter systems. It leads to a wealth of novel quantum phenomena, including the non-Hermitian skin effect Zhao et al. (2025a); Yoshida et al. (2024); Song et al. (2019a); Okuma et al. (2020); Kawabata et al. (2023); Li et al. (2020a, b); Manna and Roy (2023); Borgnia et al. (2020), non-Hermitian topology Ochkan et al. (2024); Yao and Wang (2018); Bergholtz et al. (2021); Yang et al. (2024); Song et al. (2019b); Hamanaka et al. (2024); Kawabata et al. (2019); Gong et al. (2018), and exceptional points Lee (2022); Ding et al. (2022); Tang et al. (2020); Miri and Alù (2019). These phenomena predominantly arise in non-interacting or weakly interacting systems, where quasi-particle descriptions are well-defined and applicable. However, for strongly correlated systems where the notion of quasi-particle fundamentally breaks down, the role of non-Hermiticity remains largely unexplored.
@@
In this letter, we uncover that the interplay between nFL and nH can generate anomalous Kondo conductance behaviors absent in all existing Kondo systems. To ground our theory in physical settings, we propose a practical setup that realizes the non-Hermitian multichannel Kondo (nHMCK) model (Eq.(1) below). The perturbative renormalization group (RG) analysis predicts both a weak-coupling and a strong-coupling phase, as shown in Fig.1. Around the weak-coupling FP, the impurity moment decouples from the conduction electrons even for antiferromagnetic Kondo couplings, leading to a universal low-temperature conductance, $\sim 1/\mathrm{ln}^{2}(T/T_{K})$. Although the conductance is similar to that of the underscreened Kondo models  Posazhennikova and Coleman (2005); Coleman and Pépin (2003); Mehta et al. (2005), it is driven by a completely different dissipation mechanism–the nH-induced decoupling. Beyond the weak-coupling phase, Bethe ansatz approach gives richer physics than that predicted by the perturbative RG analysis Kattel et al. (2025a, b), which is verified by our non-Hermitian numerical renormalization group (NRG) approach. Importantly, after imposing $\mathcal{PT}$ symmetry, it is found that the low-energy physics of the strong coupling phase behaves similarly with that of the Hermitian case, and is described by a boundary conformal field theory (BCFT). Remarkably, the fractionalized nature of the nFL, when combined with non-Hermiticity, causes an anomalous deviation in the Kondo conductance with raising the temperature (Fig.4 below). Such a nH-enriched anomaly stems from an emergent BCFT with non-Hermitian boundary operators, which is unique to systems exhibiting both nH and nFL. These results point to a novel class of transport phenomena unrecognized before, driven simultaneously by correlation and dissipation.
@@
Away from the weak-coupling FP, the situation is more subtle, as the perturbative analysis could breakdown. Along the red arrow in Fig.3(b), a critical curve (marked by yellow) occurs, separating the weak and strong non-Hermiticity region. Notably, for an intermediate non-Hermiticity region, Bethe ansatz studies on the single-channel case revealed an interesting Yu-Shiba-Rusinov (YSR)-like phase Kattel et al. (2025a). Here, we solve the Bethe ansatz equation for the multichannel model without $\mathcal{PT}$ symmetry, and find that the YSR-like phase still exists in this region (Sec.V of Ref. Yi et al. (2026).). In the weak non-Hermiticity region with small $\theta$, although the perturbative RG predicts the flow towards $(g_{r},g_{i})=(2/n,0)$, non-perturbative calculations are indispensable to fully understand the low-energy Kondo physics. Thus, we further utilize the non-Hermitian NRG Burke and Mitchell (2025) to solve Eq.(1), taking the $n=2$ case as an example. The details of our NRG method are presented in Ref. Yi et al. (2026).
As shown by Fig.3(e)(f), the real spectrum at large RG step coincides with that of the Hermitian case for small $\theta$, while significant deviations are observed for large $\theta$. Meanwhile, the impurity entropy saturates to $\mathrm{ln}2/2$ and $\mathrm{ln}2$ for small and large $\theta$, respectively (Fig.3(d)). These indicate that the system indeed flows to the overscreened (unscreened) phase under weak (strong) non-Hermiticity. For small $\theta$, the entropy decreases monotonously with temperature (blue curve in Fig.3(d)), displaying the standard Kondo regime behavior. For large $\theta$, the local moment phase is identified for both the high and low temperature regime, again validating the circling RG flow in Fig.3(b)  Nakagawa et al. (2018). Remarkably, for both large and intermediate $\theta$, the impurity entropy exhibits a non-monotonous evolution and can exceed the free local moment entropy, $\ln 2$, in an intermediate temperature range (red curve in Fig.3(d)). Such anomalous behaviors were also reported in superconductors with an impurity  Kattel et al. (2025c, d) . They reflect the fact that additional degrees of freedom emerge and contribute to the impurity entropy, providing clear evidence of the YSR phase predicted by Ref. Kattel et al. (2025a).
@@
Conclusion and discussion.– We have shown that non-Hermiticity can generate unconventional Kondo conductance behaviors. Particularly, when nFL meets nH, an anomalous upturn would occur, originating from an emergent BCFT with non-Hermitian boundary operators. The non-Hermitian LIO identified here is unique for Kondo systems that simultaneously display nH and nFL. For example, for the non-Hermitian Fermi liquid FP (which occurs for $n=1$) , the impurity is exactly screened, and the LIOs around the strong-coupling FP are simply four-fermion operators without imaginary couplings  [^3] The four-fermion operators with imaginary couplings break $\mathcal{PT}$, thus are prohibited around the FP.. Here, we emphasize that these BCFT results are applicable only to $\mathcal{PT}$ symmetric model, while for the asymmetric cases, our NRG calculations have confirmed the emerging YSR phase Kattel et al. (2025a, b), beyond BCFT descriptions. In addition, it is worthwhile to consider generalizations to other correlated systems exhibiting nFL properties, such as the Luttinger liquid Lee and Toner (1992); Fröjdh and Johannesson (1995); Furusaki and Nagaosa (1994); Han et al. (2023); Tang et al. (2024). Furthermore, we remark that multi-orbital cold-atom systems also provide a feasible platform to realize our predicted Kondo phenomena, as they can be engineered to host both localized and itinerant degrees of freedom  Gorshkov et al. (2010); Riegger et al. (2018). Designable dissipation and system-integrable monitoring equipments are also achievable in this platform  Syassen et al. (2008); Takasu et al. (2020); Bouganne et al. (2020); Zhao et al. (2025b, c); Vanhoecke and Schirò (2025). Last, our predicted anomalous conductance could serve as an additional signature of Majorana fermions. Given its strong-correlation and nonlocal origin, this signature may help distinguish Majorana modes from trivial bound states in experimental measurements.
@@ Acknowledgements.
W. Z. Y and Y. C contributed equally to this work. We acknowledge Wei-Qiang Chen for inspiring discussions. This work was supported by the National Natural Science Foundation of China (No.12322402, No.12274206, No.125B2076), the Scientific Research Innovation Capability Support Project for Young Faculty (SRICSPYF-ZY2025164), the National R&D Program of China (2024YFA1410500, 2022YFA1403601), the Quantum Science and Technology-National Science and Technology Major Project (Grant No.2021ZD0302800), the Natural Science Foundation of Jiangsu Province (No.BK20233001), and the Fundamental Research Funds for the Central Universities (KG202501).